Thursday, July 28, 2011
Today, Ben and I finished working on our slides for our presentation to cap off the research. Based on suggestions from Prof Andresen and John, we made changes to our data arrangement and delivery. We did a practice session with John spotting out more ways to improve the presentation. We were finally ready but the end of the day and gave our presentation. I think it went pretty well and we are now assured that the rest of the Physics research have a good idea of what we did during the summer!
Tuesday, July 26, 2011
We got the results today and it was another disappointment. The Na concentration was fluctuating while the Mg concentration seemed better but not still up to expectation. Professor Andresen decided to check the parts of the spectrometer again to see if there is any problem wrong with it and he also personally made some solutions to see what happens. Ben and I also started working on our slide show for the presentation we have tomorrow. We are gathering our data and doing some final analysis on them.
Monday, July 25, 2011
I remade the solutions using the right concentrations of Mg today. The results we great! What a sense of relief. All the prepared concentrations fell with 1% of their intended concentration. So we decided to step it up one notch. Professor Andresen suggested that I made more solutions in ppb But this time I should make two sets: One with only Mg diluted with water and then the other one with the complete components ( tris and Nacl). I made the solutions and stored them in the fridge so we could test them on Monday and analyze the results.
Today, we ran and analyzed the results of the solutions I made in ppb (parts per billion). There was another problem today as the results showed that there was no Mg in the solutions at all. I checked through my calculations and found out that my calculations were off by a decimal place which made my calculations for the Mg concentrations 1/10th of the actual values they should have been thereby making them undetectable by the spectrometer. I decided to repeat making the solutions tomorrow.
Today, we got the results for the solutions I made yesterday. We had both good and bad news. Good news first: The Na concentrations we almost on points as we had only two samples falling of the 10mM concentration it was supposed to be(7.9 & 7.7mM).The rest were about within 5% of the 10mM intended concentration. However, the bad news is that the Mg concentrations were all off by a factor of 7. This was surprising and confusing at the same time because I was pretty sure I did not multiply my calculations by any factor at all. So, the conclusion today was that I should make some solutions of just different Mg concentrations and water without the tris and NaCl. I made about 5 samples starting form an Mg concentration of 300ppb with an increase of 600ppb between solutions.
Today, I took special care in making the solutions. I weighed them instead of the normal approach of just pippetting the needed volumes into test tubes. I also talked to Professor Andresen about possible problems with the whole experiment and we decided that we wait for the results of this run and then we can change our approach to making the buffer solutions.
Tuesday, July 19, 2011
After the little success we got, I decided to make more solutions today. However, the results are not looking good. The next step will have to be for me to weigh them instead of just relying on the pipette to measure the amount of solution that I put into the test tubes.
Monday, July 18, 2011
Today I ran two large sets of samples, one for myself and one for Fash. Fash's samples were his latest set of buffers, for which we have high hopes. After fixing our problems last week, these results should be very consistent. We will look at the concentration data tomorrow. I finished running all of my Cohex 10mM series, and the preliminary results show a clear pattern. As expected, at high Mg concentrations, very little DNA precipitated, so the results are somewhat unreliable. Tomorrow I will double check the buffer solution, and begin to write up some summaries of our data.
Posted by Anonymous at 6:55 PM
Last week was a busy week, which kept me from posting much. I'll try to give a somewhat decent account of what happened, and the status of our projects. Last week, we had some trouble getting consistent results from Fash's buffers. We investigated a variety of variables, from my calibration technique to pippetting errors. We also had a problem with counts from the spectrometer. We eventually discovered that part of our problem may have been coming from the plasma torch area being dirty. There was also a possible problem with consistent nitrogen flow. We now think we have most of the issues figured out, and should be able to proceed with collecting consistent results. Ben has been getting consistent buffer values, and is proceeding with his NCP experiments. For my project, last week I ran the second buffer solution, and began to test the actual samples. The second buffer solution had somewhat surprising concentrations, because they came out to consistently around 8 mM Cohex, when we thought they were 10. I am going to double check these results after I test the actual samples further.
Posted by Anonymous at 6:41 PM
Today we tested johns results, and although they looked better than mine, they are not as close as we wanted them to be. So we were left wondering what was wrong with either the machine, the stocks or us! Well, we ruled the machine out because Ben's results were good. We were on the verge of ordering new stock solutions when professor Andresen decided to check the sprayer in the machine( spectrometer) and found out it was kind of clogged and worn out. He changed it with a new one and we hoped that as the problem with the machine. So I made some more solutions are we quickly tested them. Surprisingly we got really good results but there was a little problem because the measured concentrations of Mg we measured were in multiples of 10 of what I actually prepared. This might have been a mistake on my part so we decided to make more solutions tomorrow.
We tested and analyzed the samples I made yesterday. They still turned out random. So John and I decided that he made a few samples of his own so we could conclude if the problem was with my measurement or the stock solutions themselves. John made solutions with magnesium concentrations of 1mM and 2mM. We will test the results tomorrow.
Tuesday, July 12, 2011
After the conference we each have some new assignments. Mine - to test the ionic attachment of Magnesium to NCP. It's simular to the last tasks, but a different method. Instead of adding Mg solutions of different concentrations to NCP, we are adding NCP to pre tested concentrations of Mg solutions, and testing the difference between the two. The goal is to get really careful and accurate data, so yesterday and today was spent working on the method. Yesterdays concentrations had a 6 sample standard divination of 1.4%, hopefully today's samples will be even less.
After a week long break, we all gathered together today to discuss. Professor Andresen gave us a little briefing of how the conference went and he proceeded to let us know about future experiments we were going to carry out. The first problem we wanted to correct was the ability to make solutions to a good accuracy. So , I made buffer solutions of different Mg concentrations. If the results do not turn out well, I will have to make some more tomorrow too because we are gunning for perfection with this.
Sunday, July 3, 2011
Today, we got the results of the last test which turned out pretty well but with some unusual uncertainties. However, the got some good results on the concentrations at which the precipitation of the NCP occurs. I also helped Professor Andresen do some spins with some new NCP samples. The only difference with this samples was that the H3 and H4 tails were cut off from them. Professor Andresen was also preparing for his conference next week.